zeeawad:

In his fish pond, Graham has goldfish and clownfish in the ratio 4:7. If he adds 6 more goldfish, he has the same number of each type of fish. How many fish did he have in the first place?

Extrinix:

So let's start off by finding the total amounts of each fish type We are giving 4:7 goldfish to clownfish, and if you add 6 goldfish, it equals the amount of clownfish, as stated, so we need to figure out what times what gives you an answer on both sides that equals the other \(7\times ? = 4\times ?+6\) This would lead you to \(2\), given that both equal to \(14\) \(7\times 2 = 4\times2+6\) Both equal \(14\) So, given that 6 fish were added to the goldfish to equal the clownfish, you subtract \(6\) from the \(14\) to get the amount of goldfish \(14-6=Goldfish\) \(8=Goldfish\) So that leaves you with \(14=Clownfish\) \(8=Goldfish\)

zeeawad:

thanks for your swift reply

Extrinix:

No problem ^^